OpenStudy (anonymous):

A bag has 1 red marble, 4 blue marbles, and 3 green marbles. Peter draws a marble randomly from the bag, replaces it, and then draws another marble randomly. What is the probability of drawing 2 blue marbles in a row? Explain your answer.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got a clue or you want to go real slow?

OpenStudy (anonymous):

hi again lol

OpenStudy (anonymous):

I kinda have a clue but im not sure

OpenStudy (misty1212):

ok so lets take it step by step how many marbles are in the bag total?

OpenStudy (anonymous):

8

OpenStudy (misty1212):

ok and how many are blue?

OpenStudy (anonymous):

4

OpenStudy (misty1212):

ok good, so what is the probability that the first one selected is blue?

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

4/8?

OpenStudy (misty1212):

make a guess, if you are wrong i will tell you

OpenStudy (misty1212):

yeah \[\frac{4}{8}\] except that you really want to say \(\frac{1}{2}\) same thing

OpenStudy (anonymous):

oh ok

OpenStudy (misty1212):

now you put the marble back (it says that) so you still have 4 blue marbles in the bag and 8 total

OpenStudy (anonymous):

oh so it would be 1/2 again?

OpenStudy (misty1212):

yes and for them both to happen, blue and then blue again, the probability is the product of those two

OpenStudy (misty1212):

i.e. \[\frac{1}{2}\times \frac{1}{2}\]

OpenStudy (misty1212):

hope that is more or less clear

OpenStudy (anonymous):

Yeah makes sense

OpenStudy (misty1212):

k good